### v1

# Operator-valued Khintchine inequality for $\epsilon$-free semicircles

Abstract.
We exhibit several bounds for operator norms of the sum of $\epsilon$-free semicircular random variables introduced in the paper of Speicher and Wysoczański. In particular, using the first and second largest eigenvalues of the adjacency matrix $\epsilon$, we show analogs of the operator-valued Khintchine-type inequality obtained by Haagerup and Pisier.
1. Introduction
The mixture of classical and free independence has been considered for many years in different contexts. In group theory, Green [Gre90] introduced the graph product of groups in 1990, which is the mixture of direct and free products associated with a given graph. After that, Caspers and Fima [CF17] imported this graph product into the field of operator algebras. Independently, Młotkowski [Mo04] studied the same object (he called it $\Lambda$-freeness) from the perspective of non-commutative probability theory, which was revisited by Speicher and Wysoczański [SW16] (they called it $\epsilon$-freeness). In recent years, there have been a couple of results found in different directions. Charlesworth et.al. [CdSH+24a, CdSH+24b] proved several fundamental properties (factoriality, amenability, fullness, etc.) of the graph product of von Neumann algebras and also constructed random permutation matrix models which asymptotically obey this graph independence. Magee and Thomas [MT23] proved strong convergence of random matrix models for the graph independence, which is applied to the fundamental group of a closed hyperbolic manifold. There is also a result by Cébron et.al. [CSY24] about the central limit theorem for $\epsilon$-freeness such that the limit distribution can be characterized by graphon.
One of our motivations is the paper by the first-named author and his coauthors [CYZZ23] where they study the sum of random matrices that are asymptotically $\epsilon$-free. In their paper, the authors consider this sum as random local Hamiltonians based on the paper by Charlesworth and the first author [CC21], which also appears in the context of quantum many-body system [ML19]. Recently, Chen et al. [CGVvH24] proved the strong convergence of their (Gaussian) random matrix model, which says the operator norm of any polynomial in them converges to the operator norm of the same polynomial in $\epsilon$-free semicircles. In our paper, we show several bounds for the operator norm of the sum of $\epsilon$-free semicircle random variables. For upper bounds, we have the following result.
Theorem 1.1.
Let $s_{1},\dots,s_{d}$ be a standard $\epsilon$-free semicircle. Then, for any $a_{1},\ldots,a_{d}\in B(H)$, we have
$\left\|\sum_{i=1}^{d}a_{i}\otimes s_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}\leq 2\sqrt{\lambda_{1}+1}\max\left\{\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\right\|_{B(H)}^{\frac{1}{2}},\ \left\|\sum_{i=1}^{d}a_{i}a_{i}^{*}\right\|_{B(H)}^{\frac{1}{2}}\right\}$
where $\lambda_{1}$ is the largest eigenvalue of the adjacency matrix of $\epsilon$. Moreover, when $\epsilon$ is a connected regular graph with a degree less than $d-1$, we also have
$\left\|\sum_{i=1}^{d}a_{i}\otimes s_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}\leq 2\sqrt{\frac{d(\lambda_{2}+1)}{d-(\lambda_{1}-\lambda_{2})}}\max\left\{\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\right\|_{B(H)}^{\frac{1}{2}},\ \left\|\sum_{i=1}^{d}a_{i}a_{i}^{*}\right\|_{B(H)}^{\frac{1}{2}}\right\}$
where $\lambda_{2}$ is the second largest eigenvalue of the adjacency matrix $\epsilon$.
This result can be seen as an analog of the operator-valued Khintchine inequality observed by Haagerup and Pisier [HP93] for the free group $\mathbb{F}_{d}$;
$\left\|\sum_{|g|=1}a_{g}\otimes\lambda(g)\right\|_{B(H\otimes l^{2}(\mathbb{F}_{d}))}\leq 2\max\left\{\left\|\sum_{|g|=1}a_{g}^{*}a_{g}\right\|_{B(H)}^{\frac{1}{2}},\ \left\|\sum_{|g|=1}a_{g}a_{g}^{*}\right\|_{B(H)}^{\frac{1}{2}}\right\}$
where $a_{g}\in B(H)$. We note that Caspers, Klisse, and Larsen [CKL21] also proved a Khintchine-type inequality for the graph product. Although they use the information of cliques (that is, complete subgraphs), our inequality uses the eigenvalues of the adjacency matrix and is also optimal in several cases (see Remark 3.3).
For lower bounds, we were unable to use the eigenvalues of the adjacency matrix $\epsilon$. Instead, we have a bound that involves the cliques in $\epsilon$, which is slightly better than the trivial bound.
Theorem 1.2.
Let $s_{1},\ldots,s_{d}$ be $\epsilon$-free semicircular elements. Then we have
[…]
where $\mathrm{Cliq}(\epsilon)$ is the set of cliques of $\epsilon$.
For the proofs, we use a representation of $\epsilon$-free semicircles on a kind of Fock spaces. This method has been applied to estimate operator norms of polynomials in $q$-Gaussian system [Boż99] and twisted-Gaussian system [Kró05]. We hope that our results will be a trigger for further studies of the analytic properties of graph-independent random variables.
Acknowledgement
We are very indebted to Liang Zhao for multiple discussions during his visit to Kyoto, and in particular for explaining $\epsilon$-freenss to the second-named author. The authors thank Roland Speicher for the useful discussions and for telling us about the unpublished version of his paper. We also thank Zhiyuan Yang for telling us about the operator-valued Khintchine-type inequality. B. C. was supported by JSPS Grant-in-Aid Scientific Research (B) no. 21H00987, and Challenging Research (Exploratory) no. 23K17299. A. M. was supported by JSPS Research Fellowships for Young Scientists, JSPS KAKENHI Grant Number JP 22J12186, JP 22KJ1817, and JSPS Overseas Research Fellowship.
2. Preliminaries
2.1. $\epsilon$-independence
Throughout the paper, $\epsilon$ denotes a simple (i.e., no multiple edges) undirected graph without self-loops. Let $d\in\mathbb{N}$ be the number of vertices in $\epsilon$, and we assign different numbers in $\{1,\ldots,d\}$ to the vertices in $\epsilon$. We also identify the graph $\epsilon$ with its adjacency matrix $\epsilon=(\epsilon_{ij})_{1\leq i,j\leq d}$, which is a symmetric matrix whose $(i,j)$ entries are $1$ if the $i$-th vertex is connected to the $j$-th vertex and 00 otherwise. Note that diagonal entries of the adjacency matrix are 00 since $\epsilon$ does not have self-loops. Based on the definition in [SW16] by Speicher and Wysoczański, the $\epsilon$-freeness is defined as follows.
Definition 2.1.
For $n\in\mathbb{N}$, let $I_{n}^{\epsilon}$ be the set of elements $(i_{1},\ldots,i_{d})\in\{1,\ldots,d\}^{n}$ such that if there exists $1\leq k<l\leq n$ with $i_{k}=i_{l}$, then there is a $k<m<l$ with $\epsilon_{i_{k},i_{m}}=0$. Let $(\mathcal{A},\phi)$ be a non-commutative probability space. Then, we say unital subalgebras $\{\mathcal{A}_{i}\}_{i=1}^{d}$ of $\mathcal{A}$ are $\epsilon$-free independent if they satisfy
  • If $\epsilon_{ij}=1$, then $\mathcal{A}_{i}$ and $\mathcal{A}_{j}$ commute.
  • For each $n\in\mathbb{N}$, whenever we take $a_{k}\in\mathcal{A}_{i_{k}}$ ($k=1,\ldots,n)$ such that $\phi(a_{k})=0$ and $(i_{1},\ldots,i_{n})\in I_{n}^{\epsilon}$, we have
    $\phi[a_{1}a_{2},\cdots a_{n}]=0.$
2.2. Representation of $\epsilon$-free semicircles
We briefly discuss a representation of $\epsilon$-free semicircular random variables, which has been known by Speicher and Wysoczański [SW16, Proposition 5.1 in the unpublished version], and Magee and Thomas [MT23, Section 2.3.]. In fact, Speicher and Wysoczański [SW16, Proposition 5.1 in the unpublished version] showed that $\epsilon$-free semicircles have the same joint (non-commutative) distributions as the mixed $(\epsilon_{ij})$-Gaussians which are represented as the sum of creation and annihilation operators $l_{i}+l_{i}^{*}$ satisfying the following mixed $(\epsilon_{ij})$-relation [Spe93, BS94];
$l_{i}^{*}l_{j}-\epsilon_{ij}l_{j}l_{i}^{*}=\delta_{ij}I.$
Remark 2.2.
In [Kró05, Theorem 29], Królak shows a Haagerup-type inequality for the mixed $(q_{i,j})$-Gaussians with $\max|q_{i,j}|<1$ that estimates the operator norm of the sum of the generators. However, we cannot directly apply this inequality to $(\epsilon_{ij})$-Gaussians since $\epsilon_{ij}$ is 00 or $1$. Thus, we need to treat $\epsilon$-free semicircles separately.
Here, we give an alternative construction of $\epsilon$-free semicircles that simplifies their manipulation. Let $\epsilon$ be given. We define an equivalence relation, denoted by $\sim_{\epsilon}$, on the set of words $[d]^{*}=\bigsqcup_{n=0}^{\infty}\{1,\ldots,d\}^{n}$ with the empty word 00 (that is, $\{1,\ldots,d\}^{0}=\{0\}$) by
$uijv\sim_{\epsilon}ujiv\quad\text{if and only if}\quad\epsilon_{ij}=1$
where $u,v\in[d]^{*}$. Let $[d]_{\epsilon}^{*}$ denote the equivalence class of $[d]^{*}$ with respect to $\sim_{\epsilon}$. Let us consider the following Hilbert space
$\mathcal{F}_{\epsilon}=\bigoplus_{w\in[d]_{\epsilon}^{*}}\mathbb{C}e_{w}$
where $\mathbb{C}e_{w}$ is a one-dimensional Hilbert space with unit vector $e_{w}$ and we see $e_{0}$ as the vacuum. As in the full Fock space, we define left creation operators $\{l_{i}\}_{i=1}^{d}$ by
$l_{i}e_{w}=e_{iw}.$
Note that $l_{i}$’s are isometries on $\mathcal{F}_{\epsilon}$ since $iu\not\sim_{\epsilon}iv$ if $u\not\sim_{\epsilon}v$. In this setting, the left annihilation operator $l_{i}^{*}$ is given by
$l_{i}^{*}e_{w}=\begin{cases}e_{w^{\prime}}\ \text{if}\ w\sim_{\epsilon}iw^{\prime}\\ 0\ \text{otherwise.}\end{cases}$
Note that since we have $u\sim_{\epsilon}v$ if $iu\sim_{\epsilon}iv$, $l_{i}^{*}$ is well-defined and satisfies
$\langle l_{i}^{*}e_{u},e_{v}\rangle=\langle e_{u},l_{i}e_{v}\rangle.$
Moreover, one can directly check the $(\epsilon_{ij})$-relation
$l_{i}^{*}l_{j}-\epsilon_{ij}l_{j}l_{i}^{*}=\delta_{ij}I.$
Put $s_{i}=l_{i}+l_{i}^{*}$. One sees that each $s_{i}$ has the standard semicircle distribution with respect to the vacuum state. This construction is a rephrasing of the construction in [Spe93, BS94] (with $q_{ij}=\epsilon_{ij}$) and [MT23, Section 2.1] using an orthonormal basis $\{e_{w}\}_{w\in[d]^{*}_{\epsilon}}$. Thus, we have the following
Theorem 2.3.
In the setting above, $\{s_{i}\}_{i=1}^{d}$ forms $\epsilon$-free independent semicircle distributions with respect to the vacuum state $\tau(\cdot)=\langle\ \cdot\ e_{0},e_{0}\rangle$.
3. Main results
Throughout this section, let $H$ be a Hilbert space and $B(H)$ be the set of bounded operators on $H$. In addition, $\mathcal{F}_{\epsilon}$ denotes the Hilbert space described in the previous section.
First of all, we prove an analog of the operator-valued Khintchine inequality observed by Haagerup and Pisier [HP93] for the free group $\mathbb{F}_{d}$;
$\left\|\sum_{|g|=1}a_{g}\otimes\lambda(g)\right\|_{B(H\otimes l^{2}(\mathbb{F}_{d}))}\leq 2\max\left\{\left\|\sum_{|g|=1}a_{g}^{*}a_{g}\right\|_{B(H)}^{\frac{1}{2}},\ \left\|\sum_{|g|=1}a_{g}a_{g}^{*}\right\|_{B(H)}^{\frac{1}{2}}\right\}$
where $a_{g}\in B(H)$.
Theorem 3.1.
Let $s_{1},\dots,s_{d}$ be standard $\epsilon$-free semicircles. Then, for any $a_{1},\ldots,a_{d}\in B(H)$, we have
$\left\|\sum_{i=1}^{d}a_{i}\otimes s_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}\leq 2\sqrt{\lambda_{1}+1}\max\left\{\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\right\|_{B(H)}^{\frac{1}{2}},\ \left\|\sum_{i=1}^{d}a_{i}a_{i}^{*}\right\|_{B(H)}^{\frac{1}{2}}\right\}$
where $\lambda_{1}$ is the largest eigenvalue of the adjacency matrix $\epsilon$. Moreover, when $\epsilon$ is a connected regular graph with a degree less than $d-1$, we also have
$\left\|\sum_{i=1}^{d}a_{i}\otimes s_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}\leq 2\sqrt{\frac{d(\lambda_{2}+1)}{d-(\lambda_{1}-\lambda_{2})}}\max\left\{\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\right\|_{B(H)}^{\frac{1}{2}},\ \left\|\sum_{i=1}^{d}a_{i}a_{i}^{*}\right\|_{B(H)}^{\frac{1}{2}}\right\}$
where $\lambda_{2}$ is the second largest eigenvalue of the adjacency matrix $\epsilon$.
Proof.
We use the representation $s_{i}=l_{i}+l_{i}^{*}$ on $\mathcal{F}_{\epsilon}$ in section 2.2 where we have $l_{i}^{*}l_{j}-\epsilon_{ij}l_{j}l_{i}^{*}=\delta_{ij}I$. Then, we have
$\left\|\sum_{i=1}^{d}a_{i}\otimes s_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}$ $\leq\left\|\sum_{i=1}^{d}a_{i}\otimes l_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}+\left\|\sum_{i=1}^{d}a_{i}\otimes l_{i}^{*}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}$
$=\left\|\sum_{i=1}^{d}a_{i}\otimes l_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}+\left\|\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}.$
If $\epsilon_{ij}=0$, then we have $l_{i}^{*}l_{j}=0$ ($i\neq j$) or $l_{i}^{*}l_{i}=I$. Thus we have
$\left(\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}^{*}\right)\left(\sum_{j=1}^{d}a_{j}\otimes l_{j}\right)$ $=\sum_{i=1}^{d}a_{i}^{*}a_{i}\otimes l_{i}^{*}l_{i}+\sum_{i,j=1}^{d}\epsilon_{ij}a_{i}^{*}a_{j}\otimes l_{i}^{*}l_{j}$
$=L^{*}(\epsilon+I_{d})L.$
where $L={}^{t}(a_{1}\otimes l_{1},\ldots a_{d}\otimes l_{d})$. Using the spectral decomposition, we can write $\epsilon+I_{d}=\sum_{i=1}^{d}(\lambda_{i}+1)P_{i}$ where $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}$ are the eigenvalues of $\epsilon$ and $P_{i}$ is the spectral projection with respect to $\lambda_{i}$. Thus, we have
$L^{*}(\epsilon+I_{d})L$ $=\sum_{i=1}^{d}(\lambda_{i}+1)L^{*}P_{i}L$
$\leq(\lambda_{1}+1)L^{*}\left(\sum_{i=1}^{d}P_{i}\right)L$
$=(\lambda_{1}+1)L^{*}L$
$=(\lambda_{1}+1)\sum_{i=1}^{d}a_{i}^{*}a_{i}\otimes I.$
By taking operator norms on both sides, we have
$\left\|\sum_{i=1}^{d}a_{i}\otimes l_{i}\right\|^{2}_{B(H\otimes\mathcal{F}_{\epsilon})}$ $\leq(\lambda_{1}+1)\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\otimes I\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}$
$=(\lambda_{1}+1)\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\right\|_{B(H)}$
By replacing $a_{i}$ by $a_{i}^{*}$, we also have
$\left\|\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}\right\|_{B(H\otimes\mathcal{F}_{\epsilon})}^{2}\leq(\lambda_{1}+1)\left\|\sum_{i=1}^{d}a_{i}a_{i}^{*}\right\|_{B(H)}.$
Moreover, when $\epsilon$ is a connected regular graph with a degree less than $d-1$, $P_{1}$ is the matrix whose entries are all $\frac{1}{d}$, and we have $L^{*}P_{1}L=\frac{1}{d}(\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}^{*})(\sum_{j=1}^{d}a_{j}\otimes l_{j})$
$\sum_{i=1}^{d}(\lambda_{i}+1)L^{*}P_{i}L$ $=(\lambda_{1}+1)L^{*}P_{1}L+\sum_{i=2}^{d}(\lambda_{i}+1)L^{*}P_{i}L$
$\leq(\lambda_{1}+1)L^{*}P_{1}L+(\lambda_{2}+1)L^{*}\left(\sum_{i=2}^{d}P_{i}\right)L$
$=(\lambda_{1}+1)L^{*}P_{1}L+(\lambda_{2}+1)L^{*}(I_{d}-P_{1})L$
$=(\lambda_{1}-\lambda_{2})L^{*}P_{1}L+(\lambda_{2}+1)\sum_{i=1}^{d}a_{i}^{*}a_{i}\otimes I$
$=\frac{\lambda_{1}-\lambda_{2}}{d}\left(\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}^{*}\right)\left(\sum_{j=1}^{d}a_{j}\otimes l_{j}\right)+(\lambda_{2}+1)\sum_{i=1}^{d}a_{i}^{*}a_{i}\otimes I$
By subtracting $(\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}^{*})(\sum_{j=1}^{d}a_{j}\otimes l_{j})$, we have
$(1-\frac{\lambda_{1}-\lambda_{2}}{d})\left(\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}^{*}\right)\left(\sum_{j=1}^{d}a_{j}\otimes l_{j}\right)\leq(\lambda_{2}+1)\sum_{i=1}^{d}a_{i}^{*}a_{i}\otimes I,$
and therefore we have
$\left\|\sum_{i=1}^{d}a_{i}\otimes l_{i}\right\|^{2}_{B(H\otimes\mathcal{F}_{\epsilon})}\leq\frac{d(\lambda_{2}+1)}{d-(\lambda_{1}-\lambda_{2})}\left\|\sum_{i=1}^{d}a_{i}^{*}a_{i}\right\|_{B(H)}.$
Similarly, by replacing $a_{i}$ by $a_{i}^{*}$, we obtain an inequality for $\sum_{i=1}^{d}a_{i}^{*}\otimes l_{i}$ and the desired inequality. ∎
Taking $a_{i}=1$ for any $i$, we can deduce the following corollary from the above theorem.
Corollary 3.2.
Let $s_{1},\dots,s_{d}$ be $\epsilon$-semicircles. Then we have
$\left\|\sum_{i=1}^{d}s_{i}\right\|\leq 2\sqrt{d(\lambda_{1}+1)}$
where $\lambda_{1}$ is the largest eigenvalue of $\epsilon$. Moreover, if we assume that $\epsilon$ is a connected regular graph with a degree less than $d-1$, then we also have
$\left\|\sum_{i=1}^{d}s_{i}\right\|\leq 2d\sqrt{\frac{\lambda_{2}+1}{d-(\lambda_{1}-\lambda_{2})}}$
where $\lambda_{2}$ is the second largest eigenvalue of the adjacency matrix of $\epsilon$.
Remark 3.3.
The first inequality in Corollary 3.2 is optimal in free ($\epsilon=0$) and classical independent ($\epsilon=E_{d}-I_{d}$ where all entries of $E_{d}$ is $1$) case. The second inequality is better than the first because we have
$d(\lambda_{1}+1)(1-\frac{\lambda_{1}-\lambda_{2}}{d})-d(\lambda_{2}+1)=(\lambda_{1}-\lambda_{2})(d-\lambda_{1}-1)\geq 0.$
Moreover, the second inequality is optimal in the case where multipartite complete regular graph like below
1234
In this case, the operator norm of the sum is $4\sqrt{2}$ because $s_{1}+s_{2}$ and $s_{3}+s_{4}$ are classically independent and each distribution is a semi-circle distribution with variance $2$ (i.e. $\|s_{1}+s_{2}\|=2\sqrt{2}$). On the other hand, the largest eigenvalue of the adjacency matrix is $\lambda_{1}=2$ and the second largest eigenvalue is $\lambda_{2}=0$. Note that the first inequality yields the bound $2\sqrt{12}=4\sqrt{3}$, but the second inequality gives us the optimal bound $4\sqrt{2}$. This correspondence still holds for $n$-partite complete regular graph with $m$ discrete components ($m=n=2$ in the above case).
Next, we turn to lower bounds.
Theorem 3.4.
Let $s_{1},\ldots,s_{d}$ be $\epsilon$-free semicircular elements. Then we have
[…]
where $\mathrm{Cliq}(\epsilon)$ is the set of cliques of $\epsilon$.
Proof.
Since a free semicircular system is the minimal case in terms of operator norm, we always have
$\left\|\sum_{i=1}^{d}s_{i}\right\|\geq 2\sqrt{d}.$
On the other hand, let $A=\{i_{1},\ldots,i_{k}\}$ be a clique of $\epsilon$. Then for $N\in\mathbb{N}$, we consider the vector in $\mathcal{F}_{\epsilon}$
$\xi_{N}=\sum_{j_{1},\ldots,j_{k}=1}^{N}e_{i_{1}^{j_{1}}\cdots i_{k}^{j_{k}}}.$
Since all summands are mutually orthogonal, the square of the norm of this vector is equal to $N^{k}$. Since $\{i_{1},\ldots,i_{k}\}$ is a clique, we have
$s_{i_{m}}\xi_{N}$ $=$ $\sum_{j_{1},\ldots,j_{k}=1}^{N}e_{i_{1}^{j_{1}}\cdots i_{m}^{j_{m}+1}\cdots i_{k}^{j_{k}}}+\sum_{j_{1},\ldots,j_{k}=1}^{N}e_{i_{1}^{j_{1}}\cdots i_{m}^{j_{m}-1}\cdots i_{k}^{j_{k}}}$
$=$ $\sum_{\begin{subarray}{c}1\leq j_{n}\leq N,\ n\neq m\\ 2\leq j_{m}\leq N-2\end{subarray}}2e_{i_{1}^{j_{1}}\cdots i_{k}^{j_{k}}}+\sum_{\begin{subarray}{c}1\leq j_{n}\leq N,\ n\neq m\\ j_{m}=0,1,N,N+1\end{subarray}}e_{i_{1}^{j_{1}}\cdots i_{k}^{j_{k}}}.$
By taking summation over $m$, we have
$\sum_{m=1}^{k}s_{i_{m}}\xi_{N}=\eta+\sum_{j_{1},\ldots,j_{k}=2}^{N-2}2ke_{i_{1}^{j_{1}}\cdots i_{k}^{j_{k}}}$
where $\eta$ is orthogonal to the second term and $\|\eta\|^{2}=O(N^{k-1})$. On the other hand, if $i\not\in A$, we have
$s_{i}\xi_{N}=\sum_{j_{1},\ldots,j_{k}=1}^{N}e_{i\cdot i_{1}^{j_{1}}\cdots i_{k}^{j_{k}}}$
which is orthogonal to $\sum_{m=1}^{k}s_{i_{m}}\xi_{N}$. Thus we have
$\left\|\sum_{i=1}^{d}s_{i}\xi_{N}\right\|^{2}=4k^{2}(N-3)^{k}+(d-k)N^{k}+O(N^{k-1}),$
and we obtain
$\left\|\sum_{i=1}^{d}s_{i}\right\|\geq\lim_{N\to\infty}\frac{\|\sum_{i=1}^{d}s_{i}\xi_{N}\|}{\|\xi_{N}\|}=\sqrt{4k^{2}+d-k}.$
Example 3.5.
Let us consider the XY-model (complement graphs of line graphs, i.e. only neighboring elements are free). We assume that $s_{1}$ and $s_{d}$ are free, so that we have a regular graph. This is a connected $(d-3)$-regular graph, and the second largest eigenvalue is close to $1$ as $d\to\infty$.
To see this, note that the eigenvalues of a cyclic permutation matrix $P\in M_{d}(\mathbb{C})$ ($Pe_{i}=e_{i+1}$ mod $d$ for a canonical basis of $\mathbb{C}^{d}$) are $\exp\left[-\frac{2\pi k}{d}\sqrt{-1}\right]$ ($k=0,\ldots,d-1$) with a corresponding eigenvector $v_{k}={}^{t}\left(1,\exp\left[\frac{2\pi k}{d}\sqrt{-1}\right],\ldots,\exp\left[\frac{2\pi k(d-1)}{d}\sqrt{-1}\right]\right)$, and $E_{d}-I_{d}-(P+P^{*})$ is what we want to compute. Thus, when $d$ is odd, we have $\lambda_{1}=d-3$ with the eigenvector $v_{0}$, $\lambda_{2k}=\lambda_{2k+1}=-1-2\cos(\frac{2\pi}{d}(\frac{d+1}{2}-k))$ with the eigenvectors $v_{\frac{d+1}{2}-k},v_{\frac{d-1}{2}+k}$ for $k=1,\ldots,\frac{d-1}{2}$. When $d$ is even, we have $\lambda_{1}=d-3$ with the eigenvector $v_{0}$, $\lambda_{2}=1$ with the eigenvector $v_{\frac{d}{2}}$, and $\lambda_{2k+1}=\lambda_{2k+2}=-1-2\cos(\frac{2\pi}{d}(\frac{d}{2}-k))$ with the eingenvectors $v_{\frac{d}{2}-k},v_{\frac{d}{2}+k}$ for $k=1,\ldots,\frac{d}{2}-1$. When $d$ is odd, Corollary 3.2 gives the bound:
$\left\|\sum_{i=1}^{d}s_{i}\right\|$ $\leq 2d\sqrt{-2\cos\left(\frac{\pi(d-1)}{d}\right)\left(d-(d-3+1+2\cos\left(\frac{\pi(d-1)}{d}\right)\right)^{-1}}$
$=2d\sqrt{-\cos\left(\frac{\pi(d-1)}{d}\right)\left(1-\cos\left(\frac{\pi(d-1)}{d}\right)\right)^{-1}}.$
When $d$ is even, Corollary 3.2 gives a bound $2d\sqrt{\frac{1+1}{d-(d-3-1)}}=\sqrt{2}d$. On the other hand, by triangular inequality, we also have
$\left\|\sum_{i=1}^{d}s_{i}\right\|\leq\|s_{1}+s_{2}\|\cdot\frac{d}{2}=2\sqrt{2}\cdot\frac{d}{2}=\sqrt{2}d.$
Regarding lower bounds for the XY-model, we can take $\{2k-1\}_{k=1}^{\lceil d/2\rceil}$ as a maximal clique and we apply Theorem 1.2 to have
$\left\|\sum_{i=1}^{d}s_{i}\right\|\geq\sqrt{4\lceil d/2\rceil^{2}+d-\lceil d/2\rceil}>2\lceil d/2\rceil.$
Note that we can easily see $2\lceil d/2\rceil$ is a lower bound since $\{s_{2k-1}\}_{k=1}^{\lceil d/2\rceil}$ is a classically independent family of semicircles, so our result is a non-trivial improvement on naive lower bounds.